### v1

# Steady-State Entanglement Generation via Casimir-Polder Interactions

Abstract
We investigate the generation of steady-state entanglement between two atoms resulting from the fluctuation-mediated Casimir-Polder (CP) interactions near a surface. Starting with an initially separable state of the atoms, we analyze the atom-atom entanglement dynamics for atoms placed at distances in the range of $\sim 25$ nm away from a planar medium, examining the effect of medium properties and geometrical configuration of the atomic dipoles. We show that perfectly conducting and superconducting surfaces yield an optimal steady-state concurrence value of approximately 0.5. Furthermore, although the generated entanglement decreases with medium losses for a metal surface, we identify an optimal distance from the metal surface that assists in the generation of entanglement by the surface. While fluctuation-mediated interactions are typically considered detrimental to the coherence of quantum systems at nanoscales, our results demonstrate a mechanism for leveraging such interactions for entanglement generation.
preprint: APS/123-QED
Introduction— Nanoscale quantum systems enable efficient and tunable light-matter interactions by confining EM fields in small mode volumes D’Amico and et al (2019); Chang et al. (2018); Wang et al. (2021); Soykal and Flatté (2010); González-Tudela et al. (2024); Fang et al. (2022). Such systems hold enormous promise in the development of photonic devices that exhibit nonlinearities at the single-photon level Chang et al. (2014a); Gullans et al. (2013); Volz et al. (2014); Skljarow et al. (2022), building efficient light-matter interfaces Goban and et al (2014); Thompson and et al (2013); Rivera and Kaminer (2020); Kockum et al. (2019); Flick et al. (2018); Forn-Díaz et al. (2019) that are essential for various quantum information processing tasks et al (2018a); Laucht and et al (2021), as well as, exploring fundamental phenomena such as ultrafast dynamics at the nanoscale Koo et al. (2024); et al (2018b).
However, as quantum systems move towards nanoscales, while the coupling between atomic systems and the relevant modes of the EM field increases, so does the detrimental interaction of atomic systems with the quantum fluctuations of the EM field Sánchez-Cánovas and Donaire (2022a). For example, fluctuation-induced Casimir-Polder (CP) interactions Casimir and Polder (1948); Dutta et al. (2024) between neutral polarizable particles and nearby surfaces Klimchitskaya et al. (2009); Sánchez-Cánovas and Donaire (2022b); Sinha et al. (2018) critically limit trapping schemes and the coherence of quantum emitters near surfaces Hümmer et al. (2019); Henkel and Wilkens (1999).
The realization of practical quantum technologies requires the production and longevity of quantum coherence and entanglement MacFarlane et al. (2003); Deutsch (2020). To this end, sophisticated quantum control methods and environmental noise engineering or noise-assisted entanglement generation schemes have been proposed Evans and et al (2018); Lukin and Hemmer (2000); Neuman et al. (2020); Laucht and et al (2009). Given that fluctuation-mediated interactions are deleterious for quantum coherence and correlations of atomic systems near surfaces Sedlacek et al. (2018); Brownnutt et al. (2015); Wang et al. (2015); Kim et al. (2015); Jamonneau and et al (2016); Yang et al. (2011); Reiche et al. (2020); Jain et al. (2024); Donaire and Lambrecht (2015); Donaire et al. (2015), engineering such fluctuation phenomena to achieve better coherence and control of nanoscale quantum systems thus becomes a vital goal in developing nanoscale quantum devices Laucht and et al (2021). Furthermore, the possibility of appropriately manipulating these fluctuation-mediated interactions and turning them into a useful tool can be thus doubly advantageous Chang et al. (2014b).
In this Letter, we propose a simple and intuitive scheme for controlling fluctuation phenomena to produce steady-state entanglement between two quantum emitters near a surface. Our approach utilizes the effective coherent coupling between the atoms in the presence of a medium that dominates their surface-modified decoherence to produce robust quantum entanglement in the steady state for critical placement of the atoms relative to the surface.
Figure 1: Schematic of the model featuring two two-level atoms near a planar medium of permittivity $\epsilon(\omega)$. The atomic dipoles ${\mathbf{d}}_{1}$ and ${\mathbf{d}}_{2}$ are oriented along $\theta$ with respect to the $x$-axis.
Model—We present a system of two emitters positioned at a distance $x$ from each other and at a distance $z$ from a planar half-space medium characterized by electric permittivity $\epsilon(\omega)$ (Fig. 1). Each emitter, denoted by $i=1,2$ for the first and second emitter respectively, is modeled as a two-level system with ground state $\left|g\right\rangle_{i}$ and excited state $\left|e\right\rangle_{i}$, with a resonant transition frequency $\omega_{0}$. In the presence of the planar surface, the electric field originating from vacuum fluctuations at position $\mathbf{r}$ is defined as Buhmann (2012):
[…]
Here, $\hat{f}_{\lambda}({\mathbf{r}},\omega)$ and $\hat{f}_{\lambda}^{\dagger}({\mathbf{r}},\omega)$ are bosonic annihilation and creation operators, representing noise polarization ($\lambda=e$) and magnetization ($\lambda=m$) excitations at frequency $\omega$. The coefficients $\bar{G}_{\lambda}({\mathbf{r}},{\mathbf{r}}^{\prime},\omega)$ are related to the Green’s tensor $\bar{G}({\mathbf{r}},{\mathbf{r}}^{\prime},\omega)$, which satisfies the inhomogeneous Helmholtz equation in the presence of the medium (see SM for details). The propagator $\bar{G}({\mathbf{r}},{\mathbf{r}}^{\prime},\omega)=\bar{G}_{\text{free}}(\mathbf{r},\mathbf{r}^{\prime},\omega)+\bar{G}_{\text{sc}}(\mathbf{r},\mathbf{r}^{\prime},\omega)$ consists of the free space and scattering components. The total Hamiltonian, denoted as $\hat{H}_{T}=\hat{H}_{S}+\hat{H}_{F}+\hat{H}_{\text{int}}$, comprises of the atomic Hamiltonian $\hat{H}_{S}=\sum_{i=1}^{2}\hbar\omega_{0}\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-}$, the surface-assisted electromagnetic field in vacuum state $\hat{H}_{F}$, and the interaction Hamiltonian. $\hat{\sigma}_{i}^{+}=(\hat{\sigma}_{i}^{-})^{\dagger}=\left|e\right\rangle_{i}\left\langle g\right|_{i}$ indicate the ladder operators for the atomic transition. The interaction Hamiltonian is described by $\hat{H}_{\text{int}}=-\sum_{i=1}^{2}\hat{d}_{i}\cdot\hat{E}(\textbf{r}_{i})$, where $\hat{d}_{i}=\textbf{d}_{i}\hat{\sigma}_{i}^{+}+\textbf{d}_{i}^{*}\hat{\sigma}_{i}^{-}$ represents the electric-dipole operator associated with the $i$-th emitter located at position $\textbf{r}_{i}$, with $\textbf{d}_{i}$ denoting the electric dipole moment.
The time evolution of the density matrix $\rho_{S}$ for the emitters is governed by the Born-Markov master equation Breuer and Petruccione (2002):
$\frac{d\hat{\rho}_{S}}{dt}=-\frac{i}{\hbar}[\hat{\mathcal{H}}_{\text{eff}},\hat{\rho}_{S}]+\mathcal{L}_{\text{eff}}[\hat{\rho}_{S}]$ (2)
Here ^$\hat{\mathcal{H}}_{\text{eff}}$ is the atomic effective Hamiltonian in interaction picture defined by Sinha et al. (2018); Olivera et al. (2022); Sone et al. (2024)
^eff=subscript^$\hat{\mathcal{H}}_{\text{eff}}=$ $\hbar\left[\sum_{i=1,2}\left(\Omega_{i}^{(+)}(\mathbf{r}_{i})\hat{\sigma}_{i}^{+}\hat{\sigma}_{i}^{-}+\Omega_{i}^{(-)}(\mathbf{r}_{i})\hat{\sigma}_{i}^{-}\hat{\sigma}_{i}^{+}\right)\right.$
$\left.+\sum_{i\neq j}\Omega_{ij}(\mathbf{r}_{i},\mathbf{r}_{j})\hat{\sigma}_{i}^{-}\hat{\sigma}_{j}^{+}\right],$ (3)
and the effective surface-modified Liouvillian in Eq. (2) is given by
$\mathcal{L}_{\text{eff}}[\hat{\rho}_{S}]$
$=\sum_{ij}\frac{\Gamma_{ij}(\mathbf{r}_{i},\mathbf{r}_{j})}{2}\left(2\hat{\sigma}_{i}^{-}\hat{\rho}_{S}\hat{\sigma}_{j}^{+}-\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}\hat{\rho}_{S}-\hat{\rho}_{S}\hat{\sigma}_{i}^{+}\hat{\sigma}_{j}^{-}\right),$ (4)
where $\Gamma_{ij}(\mathbf{r}_{i},\mathbf{r}_{j})=\Gamma_{ij}^{\text{free}}(\mathbf{r}_{i},\mathbf{r}_{j})+\Gamma_{ij}^{\text{sc}}(\mathbf{r}_{i},\mathbf{r}_{j})$ corresponds to the dissipative coupling coefficient between two atoms, while $\Gamma_{ii}$ represents the spontaneous emission rate for the excited state of the $i$th atom.
The diagonal elements of the effective Hamiltonian ^$\hat{\mathcal{H}}_{\text{eff}}$ are the CP shift contributions defined as:
$\Omega_{i}^{(-)}(\mathbf{r}_{i})$ $=\frac{\mu_{0}\omega_{0}}{\pi}\int_{0}^{\infty}\frac{d\zeta~{}\zeta^{2}}{\zeta^{2}+\omega^{2}}\mathbf{d}_{i}^{*}\cdot[\bar{G}_{\mathrm{\text{sc}}}(\mathbf{r}_{i},\mathbf{r}_{i},i\zeta)]\cdot\mathbf{d}_{i}$
$\Omega_{i}^{(+)}(\mathbf{r}_{i})$ $=-\Omega_{i}^{(-)}(\mathbf{r}_{i})+\Omega_{i}^{\mathrm{\text{res}}}(\mathbf{r}_{i}),$ (5)
where, the free component of Green’s tensor is neglected, assuming it has already been taken into account as a contribution to Lamb shifts in the bare levels. Additionally, the resonant contribution $\Omega_{i}^{\mathrm{\text{res}}}(\mathbf{r}_{i})=-\mu_{0}\omega_{0}^{2}\mathrm{Re}[\mathbf{d}_{i}^{*}\cdot\bar{G}_{\mathrm{\text{sc}}}(\mathbf{r}_{i},\mathbf{r}_{i},\omega_{0})\cdot\mathbf{d}_{i}]$, is dependent exclusively on the environment’s response at the resonant frequency where $\mu_{0}$ is the Permeability of free space.
The off-diagonal elements $(i\neq j)$ of the effective Hamiltonian, along with the dissipative coupling coefficients, are determined by the resonant contribution of the real and imaginary parts of Green’s function. This is expressed as $\Omega_{ij}(\mathbf{r}_{i},\mathbf{r}j)=\Omega{ij}^{\mathrm{\text{free}}}(\mathbf{r}_{i},\mathbf{r}j)+\Omega{ij}^{\mathrm{\text{sc}}}(\mathbf{r}_{i},\mathbf{r}_{j})$ as follows:
$\Omega_{ij}^{\mathrm{\text{free},\text{sc}}}(\mathbf{r}_{i},\mathbf{r}_{j})=$ $-\mu_{0}\omega_{0}^{2}\mathrm{Re}[\mathbf{d}_{i}^{*}\cdot\bar{G}_{\mathrm{\text{free},\text{sc}}}(\mathbf{r}_{i},\mathbf{r}_{j},\omega_{0})\cdot\mathbf{d}_{j}],$
$\Gamma_{ij}^{\mathrm{\text{free},\text{sc}}}(\mathbf{r}_{i},\mathbf{r}_{j})=$ $2\mu_{0}\omega_{0}^{2}~{}\mathrm{Im}[\mathbf{d}_{i}^{*}\cdot\bar{G}_{\mathrm{\text{free},\text{sc}}}(\mathbf{r}_{i},\mathbf{r}_{j},\omega_{0})\cdot\mathbf{d}_{j}].$ (6)
Results—We analyze the dynamics of the emitters, governed by the master equation (2). As illustrated in Fig. 1, we specify $x_{1}-x_{2}\equiv x$ and note that the two atoms are positioned at the same distance $z$ from the surface. Consequently, in accordance with Eqs. (Steady-State Entanglement Generation via Casimir-Polder Interactions) and (Steady-State Entanglement Generation via Casimir-Polder Interactions), we can simplify the system by setting $\Omega_{1}^{(\mp)}(z)=\Omega_{2}^{(\mp)}(z)\equiv\Omega^{(\mp)}$, $\Omega_{12}^{(\mp)}(x,z)=\Omega_{21}^{(\mp)}(x,z)\equiv\Omega_{12}^{(\mp)}$, $\Gamma_{11}(z)=\Gamma_{22}(z)\equiv\Gamma$, and $\Gamma_{12}(x,z)=\Gamma_{21}(x,z)\equiv\Gamma_{12}$.
We introduce the atomic density matrix as $\hat{\rho}_{S}=\rho_{11}\left|ee\right\rangle\left\langle ee\right|+\rho_{22}\left|eg\right\rangle\left\langle eg\right|+\rho_{33}\left|ge\right\rangle\left\langle ge\right|+\rho_{23}\left|eg\right\rangle\left\langle ge\right|+\rho_{32}\left|ge\right\rangle\left\langle eg\right|+\rho_{44}\left|gg\right\rangle\left\langle gg\right|$. Considering the initial state of the two atoms as $\rho_{S}(t_{0})=\left|eg\right\rangle\left\langle eg\right|$, the time-dependent elements of the density matrix can be obtained from solving the master equation (for more details, see SM )
[…]
$\rho_{\text{steady}}=\frac{1}{2}\left(\left|\psi_{sub}\right\rangle\left\langle\psi_{sub}\right|+\left|gg\right\rangle\left\langle gg\right|\right),$ (8)
where $\left|\psi_{\text{{sub}}}\right\rangle\equiv\left(\left|eg\right\rangle-\left|ge\right\rangle\right)/\sqrt{2}$ corresponds to the subradiant state of the emitters that is decoupled from the EM field Dicke (1954). Thus we observe that quantum correlations emerge between the two emitters, transforming the initially uncorrelated state into an entangled steady state, with a concurrence of $C(\rho_{\mathrm{steady}})=0.5$ Horodecki et al. (2009).
The condition $\Gamma=\Gamma_{12}$ that yields the entangled steady-state signifies a balance between the emitters’ spontaneous emission rate ($\Gamma=\Gamma_{11}=\Gamma_{22}$) and the surface-mediated dipole-dipole dissipative interaction ($\Gamma_{12}$). Specifically, initiating the system with one emitter in the excited state allows for the potential delocalization of a photon between the two emitters. To elaborate further, we introduce the relative decay $D(\mathbf{r}_{1},\mathbf{r}_{2})=\Gamma-\Gamma_{12}$, which is intricately dependent on the spatial arrangement of two emitters, their dipole orientation, and the properties of the surface.
The steady-state entanglement condition thus corresponds to $D(\mathbf{r}_{1},\mathbf{r}_{2})$ approaching 0. We analyze two distinct dipole configurations, namely $zz$ and $xx$ (where atomic dipoles are perpendicular ($\theta=\pi/2$) and parallel ($\theta=0$) to the surface, respectively (see Fig. 1)), and investigate the behavior of $D(\mathbf{r}_{1},\mathbf{r}_{2})$ as a function of the emitters’ position.
The relative decay can be expressed as a sum of contributions from the free and scattering components: $D(\tilde{x},\tilde{z})=D^{\text{free}}(\tilde{x})+D^{\text{sc}}(\tilde{x},\tilde{z})$ where the scaled parameters are introduced as $\tilde{x}\equiv k_{0}x$ and $\tilde{z}\equiv k_{0}z$, with $k_{0}=\omega_{0}/c$. The free part characterizes the system in the absence of the surface, while the scattering part accounts particularly for the surface effects. Utilizing Eq. (Steady-State Entanglement Generation via Casimir-Polder Interactions), the free part is calculated as $D^{\text{free}}(\tilde{x})=\Gamma_{0}-\Gamma_{12}^{\text{free}}(\tilde{x})$. We note that the free part of the relative decay for both configurations as a function of $\tilde{x}$ starts from zero and tends to $\Gamma_{0}$ for long distances, where the dipole-dipole interaction $\Gamma_{12}^{\mathrm{free}}$ vanishes as the dipoles are infinitely separated from each other (See SM ). The scattering part of the relative decay, besides $\tilde{x}$ and dipole orientations, relies on $\tilde{z}$ and the medium properties, defined as $D^{\text{sc}}(\tilde{x},\tilde{z})=\Gamma^{sc}(\tilde{x},\tilde{z})-\Gamma_{12}^{sc}(\tilde{x},\tilde{z})$. First, we consider a perfect conductor surface with permittivity $\epsilon(\omega)\rightarrow\infty$, which is calculated for each dipole configuration based on Eq. (Steady-State Entanglement Generation via Casimir-Polder Interactions). According to calculations, as $\tilde{x}$ increases, the free and scattering components of the relative decay amplify each other in the $zz$ configuration, leading to $D_{zz}\rightarrow 2\Gamma_{0}$, while in the $xx$ configuration, their interaction is destructive, resulting in $D_{xx}\rightarrow 0$. The relative decay as a function of $\tilde{x}$ is shown in Fig. 2 (a) for each dipole configuration: in free space, denoted by $D_{zz}^{\mathrm{free}}(\tilde{x})$ and $D_{xx}^{\mathrm{free}}(\tilde{x})$, and in the presence of a perfect conductor surface, considering the constant distance $\tilde{z}=0.2$ from the surface, denoted by $D_{zz}^{\mathrm{per}}(\tilde{x},\tilde{z}=0.2)$ and $D_{xx}^{\mathrm{per}}(\tilde{x},\tilde{z}=0.2)$. It illustrates how the presence of the surface can assist with maintaining the relative decay close to zero when the dipoles are oriented parallel to the surface in the $xx$ configuration. In the inset of Fig. 3 (a), we plot the time evolution of the concurrence between two emitters at the specific point $(\tilde{x}=1,\tilde{z}=0.2)$, marked with the vertical dashed line in Fig. 3 (a), calculated based on the master equation (2). The emitters are assumed to be silicon-vacancy (SiV) centers embedded in a bulk diamond, placed near a surface with a transition wavelength $\lambda=2\pi/k_{0}=737$ nm Zhou et al. (2017). It indicates that the generated entanglement for the $xx$ configuration, with $D_{xx}^{\text{per}}\approx 0$ (solid red curve), is preserved for a long time, while for the $zz$ configuration, with $D_{zz}^{\text{per}}>0$ (dash-dotted blue curve), it decays to zero significantly faster.
The interaction between emitters and the surface can be elucidated by the dipole-image model, where the perfect conductor surface acts as a mirror for the dipoles Buhmann (2012), as depicted in Fig. 2(b-c).
Figure 2: (a) The relative decay in free space (green dashed and black dotted curves) and the total relative decay near the perfect conductor at a distance of $\tilde{z}=0.2$, plotted as functions of $\tilde{x}$ for the $zz$ (dash-dotted blue curve) and $xx$ (solid red curve) configurations. The inset shows the time-dependent concurrence of the density matrix for the system with $zz$ (dash-dotted blue curve) and $xx$ (solid red curve) configurations near the perfect conductor surface. (b) and (c), The schematic depicts the dipole-image model for the perpendicular dipoles in $zz$ configuration and the parallel dipoles in $xx$ configuration, along with their corresponding images in the medium.
Figure 3: Contour plot depicting the relative decay as a function of $\tilde{x}$ and $\tilde{z}$ in the vicinity of (a) superconducting, and (b) a gold surface. Zero relative decay is required for steady-state entanglement generation. (c) However, atom-atom entanglement can persist for a long time even if the relative decay does not vanish completely. The evolution of concurrence between atoms over time is depicted for the $xx$ configuration separated by $\tilde{x}=1$ in free space (black dotted), and at a distance $\tilde{z}=0.2$ from the perfect conductor (solid red), the superconducting (blue dash-dotted), and the gold surface (yellow dashed). The selected position ($\tilde{x}=1,\tilde{z}=0.2$) is indicated by the white $\otimes$ sign in contour plots.
In the case of $zz$ configuration, where dipoles are perpendicular to the surface, the fields of each dipole and its image reinforce each other. This reinforcement leads to a decay rate that approaches $2\Gamma_{0}$. Conversely, in the $xx$ configuration, where dipoles are oriented parallel to the surface, the dipole and its image fields tend to cancel each other out. Consequently, the emitters experience a nearly zero-field environment, resulting in minimal decay. Furthermore, in addition to the dipole-dipole interaction, each emitter is influenced by the other’s image. These interactions result in a condition where the $xx$ dipole configuration experiences almost no decay, highlighting how the surface effectively maintains the steady-state entanglement condition by minimizing the relative decay.
Focusing on $xx$ configuration, we extend our analysis to a metal surface described by the Drude model and a superconducting surface. In this representation, we model the metal surface as gold, setting the plasma frequency to $\omega_{p}\approx 1.37\times 10^{16}$ Hz and the loss parameter to $\gamma\approx 5.31\times 10^{13}$ Hz Pirozhenko et al. (2006) where the permittivity described as $\epsilon_{d}(\omega)=1-\omega_{p}^{2}/(\omega^{2}+i\omega\gamma)$ Jackson (1999). Moreover, the permittivity for a London superconductor is given by Skagerstam et al. (2006)
$\epsilon_{s}(\omega)=1-\frac{c^{2}}{\omega^{2}~{}\lambda_{L}^{2}(T)}+i~{}\frac{2~{}c^{2}}{\omega^{2}~{}\delta_{L}^{2}(T)}$ (9)
The (square of the) London penetration length and the skin depth are defined as $\lambda_{L}^{2}(T)=\lambda_{L}^{2}(0)/[1-(T/T_{c})^{4}]$, and $\delta_{L}^{2}(T)=2/(\omega\mu_{0}\sigma(T/T_{c})^{4})$, where $\sigma=2\times 10^{9}~{}\Omega^{-1}$ represents the electrical conductivity. We consider a superconducting Niobium with the critical temperature $T_{c}=8.31$ K and $\lambda_{L}(0)=35$ nm at a finite temperature $T=0.01T_{c}$.
Fig. 3 (a-b) displays contour plots of $D(x,z)$ near (a) a Niobium superconductor, and (c) a gold surface, as functions of $\tilde{x}$ and $\tilde{z}$. The transition from dark blue to red indicates an increase in relative decay from zero, with dark blue areas highlighting the parameter space where the steady-state or near steady-state entanglement condition is satisfied. According to Fig. 3(a), the relative decay tends to zero as the distance $\tilde{z}$ from the surface decreases. For the superconducting surface in the non-retarded regime, $D_{xx}^{\text{Sup}}\propto(T/T_{c})^{4}$ SM . Therefore, at low temperatures $T\ll T_{c}$, the relative decay is similar to that of a perfect conductor. In contrast, Fig. 3(b) shows that the relative decay increases as $\tilde{z}$ decreases near the gold surface. In the non-retarded regime ($\tilde{z}\ll 1$), it can be shown that $D_{xx}^{\text{Gold, sc}}(\tilde{x},\tilde{z})\propto\gamma\Gamma_{0}/\tilde{z}^{3}$. This demonstrates that in metals with non-zero loss, the emitters experience a noisier environment in the non-retarded regime as the distance from the metal surface decreases due to significant surface scattering. As illustrated in Fig. 3(b), an optimal distance from the surface can be determined to minimize the relative decay and maximize surface-mediated entanglement.
After analytically demonstrating the necessity of zero relative decay for steady-state entanglement generation, we numerically identified potential geometric configurations satisfying this condition in Fig. 3 (a)-(b). However, achieving exact zero relative decay may not be feasible in practice. Nonetheless, a significant amount of entanglement could still be maintained for an extended period between the emitters even if the relative decay does not vanish completely. According to our calculations, the relative decay of the system near a perfect conductor behaves similarly to that of the system near a superconducting surface (see SM ). The relative decay for the superconducting surface (Fig. 3 (a)) can approach zero at $\tilde{z}\leq 0.2$, resulting in nearly steady-state entanglement, which can be further enhanced as $\tilde{z}$ decreases. On the other hand, near the metal surface, the generated entanglement at $\tilde{z}<0.2$ decays quickly due to near-field interactions. However, there is an optimal distance that minimizes the relative decay. For instance, the minimum value of the relative decay at $\tilde{x}=1$ occurs at a distance of approximately $\tilde{z}\approx 0.4$ from the gold surface, resulting in maximum entanglement (Fig. 3 (b)). To compare different surfaces, we consider $\tilde{x}=1$ ($x\approx 117$ nm) and $\tilde{z}=0.2$ ($z\approx 23$ nm) to examine the entanglement generation. The selected coordinate is marked in the contour plots (Fig. 3 (a)-(b)). Furthermore, we consider a system of two emitters in free space at a distance of $\tilde{x}=1$ without the presence of a surface, which serves as a baseline for free space dipole-dipole entanglement generation. This allows us to separate and elucidate the effect of each surface.
At this point, the relative decay for each type of the surface is obtained as $D_{xx}^{\text{per}}(1,0.2)=2.2\times 10^{-3}$, $D_{xx}^{\text{Sup}}(1,0.2)=4.7\times 10^{-3}$, and $D_{xx}^{\text{Gold}}(1,0.2)=0.037$. The time evolution of the concurrence $(C(\rho_{S}(t)))$ between two emitters at this point, calculated based on the master equation (2), is plotted in Fig. 3 (c). The calculations reveal that at scaled time $\Gamma_{0}t=30$, the generated entanglement near the perfect conductor and superconducting surfaces is approximately $C^{\text{per}}_{xx}\approx 0.41$ and $C^{Sup}_{xx}\approx 0.37$, respectively, while it decays to $C^{\text{Gold}}_{xx}\approx 0.05$ near the gold surface, and for free space $C_{xx}^{\text{free}}\approx 0.1$. According to Fig. 3 (c), the concurrence is enhanced by the presence of the perfect conductor (solid red curve) and superconducting surface (dashed-dotted blue curve) compared to free space (dotted black curve), while the gold surface (dashed yellow curve) reduces entanglement generation. However, it can be shown that at point $\tilde{z}=0.4$, as an optimum distance from the surface, with the same $\tilde{x}=1$ the gold surface can also contribute positively, resulting in $C^{\text{Gold}}_{xx}\approx 0.24>C_{xx}^{\text{free}}$.
Conclusion – In conclusion, this letter introduced a method for generating steady-state entanglement by investigating the dynamics of a system with two emitters near a surface, considering the CP interaction. We analyzed the emergence of steady-state entanglement through an analytical solution of the master equation, emphasizing the necessary conditions for its realization. We introduced relative decay function $D(\mathbf{r}_{1},\mathbf{r}_{2})=\Gamma-\Gamma_{12}$ defined by spontaneous emission rates ($\Gamma\equiv\Gamma_{11}=\Gamma_{22}$) and surface-mediated dipole-dipole dissipative interaction ($\Gamma_{12}=\Gamma_{21}$). We demonstrated that achieving an entangled steady-state with a concurrence of $C(\rho_{\text{steady}})=0.5$ is possible when the relative decay, expressed as a function of the spatial positions of the emitters, their dipole orientations, and surface properties, equals zero ($D(\mathbf{r}_{1},\mathbf{r}_{2})=0$). First, we revealed that to satisfy this condition in a surface-mediated process, the emitters’ dipoles must be oriented parallel to the surface. Moreover, the relative decay as a function of the emitters’ position is affected by the surface’s optical properties. We demonstrated that the superconducting surface closely corresponds to the perfect conductor, maintaining the relative decay near zero. In contrast, due to the metal’s loss parameter, the relative decay increases near a metal surface in the non-retarded limit, leading to a faster decay of the generated entanglement. However, we identified an optimal distance from the metal surface that enhances entanglement generation compared to free space and preserves it for a longer time.
These findings have improved our understanding of entanglement dynamics in confined geometries and hold practical implications for quantum information processing protocols in nanoscale systems. The proposed method also hints at potential applications in near-surface quantum metrology and sensing, where entanglement could play a role in enhancing precision measurements and sensing capabilities.
Acknowledgments— This work was supported by the Scientific and Technological Research Council of Türkiye (TÜBİTAK) under Project Numbers 120F089 and 123F150. M. I., O. P., and Ö. M. thank TÜBİTAK for their support. K.S. acknowledges support from the National Science Foundation under Grant No. PHY-2418249, and by the John Templeton Foundation under Award No. 62422.